OpenStudy (anonymous):

PLEASE HELP ME Solve, using the substitution method. j + k = 3 j – k = 7 The solution is (8, 1) The solution is (5, –2) There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

B. The solution is (5, 2)

jimthompson5910 (jim_thompson5910):

solve j – k = 7 for j j – k = 7 j – k+k = 7+k j = 7+k

jimthompson5910 (jim_thompson5910):

then plug j = 7+k into the first equation j + k = 3 j + k = 3 7+k + k = 3 ... notice how i replaced j with 7+k solve for k

OpenStudy (anonymous):

checking my answers _____________________________

OpenStudy (anonymous):

@Haichi that's not exactly helping things. Jim was doing a great job explaining and you waltz in with an unnecessary comment.

OpenStudy (anonymous):

@tom982 thank you

OpenStudy (anonymous):

come on now, lets not get in a fight at 11pm

OpenStudy (haichi):

ok soz

OpenStudy (haichi):

Jeez

OpenStudy (anonymous):

im sorry i should have never addded on to it can we please continue

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (haichi):

not rlly

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @Haichi figure them out ur smart enough \(\color{blue}{\text{End of Quote}}\) Says the person who didn't help...

OpenStudy (anonymous):

you know what nevermind

TheSmartOne (thesmartone):

I like this quote button... :P http://prntscr.com/5cc03x

OpenStudy (anonymous):

11pm!? It's 4am here haha

jimthompson5910 (jim_thompson5910):

are you able to solve 7+k + k = 3 for k?

TheSmartOne (thesmartone):

If you have anymore questions then close this and post a new one :D

OpenStudy (anonymous):

im in florida

OpenStudy (haichi):

@TheSmartOne I helped in other ones that were so easy

OpenStudy (anonymous):

@TheSmartOne thanks idk i dont want to start a riot

OpenStudy (haichi):

U could have done them in ur sleep @TheSmartOne

OpenStudy (anonymous):

or even worst world war z

OpenStudy (dtan5457):

@Haichi No specific grade level for this section. Help and move on

TheSmartOne (thesmartone):

lol I'm leaving. So what if I could @Haichi A peron needs help, you help them you do not need to make things bigger.

OpenStudy (dtan5457):

This question is not a joke question...some people find it harder than others..

OpenStudy (haichi):

@dtan5457 No der

OpenStudy (haichi):

k @TheSmartOne

OpenStudy (anonymous):

so its B?

OpenStudy (dtan5457):

Ask Haichi he'll help you

TheSmartOne (thesmartone):

Yes it is B...

OpenStudy (anonymous):

im so smart im 12 yrs old

OpenStudy (anonymous):

B thanks

OpenStudy (anonymous):

lol jk feel free to reward me

TheSmartOne (thesmartone):

@Cholo123 You know you have to be atleast 13 years or older to be on OS right? Otherwise OS could be fined $16,000

OpenStudy (anonymous):

lol *clap emoji* @Cholo123 :)

OpenStudy (anonymous):

@TheSmartOne b for sure?

OpenStudy (anonymous):

oh i am 12 almost 13 b 13 in February. lol @welligya...

OpenStudy (anonymous):

how old r u @welligya...

OpenStudy (anonymous):

15

TheSmartOne (thesmartone):

Yes it is B for sure..

OpenStudy (anonymous):

but i have alot of work to do bye :)

OpenStudy (anonymous):

:( a boy can dream lol

OpenStudy (anonymous):

thanks

TheSmartOne (thesmartone):

No problem.

OpenStudy (abhisar):

Do not interrupt when someone is trying to teach someone.

TheSmartOne (thesmartone):

^^ Please this is not a dating website.

OpenStudy (anonymous):

:(

OpenStudy (haichi):

LOL